# A Non-Traditional Approach to Assisting Data Address Translation

### v1

@@ Abstract
This paper proposes a novel way to assist conventional data address translation. The approach, PC-Indexed Data Address Translation (PCAX), uses the PC of a load instruction, and not a data virtual address, to obtain the page table entry (PTE) for the data accessed by a load instruction. PCAX is intended to be used for a small subset of the static loads in a program. We observe that: (i) a small subset of static loads is responsible for most of the misses in a data translation lookaside buffer (DTLB), and (ii) often a dynamic instance of a static load instruction accesses the same PTE as the last dynamic instance, and consider PCAX for this subset. With PCAX the effective miss rate of a conventional DTLB can be cut down by a factor of 2-3X in many cases, and even more in some cases. PCAX is also beneficial in reducing the number of secondary TLB (STLB) misses. Since the tables used for PCAX can be accessed alongside instruction fetch, they can be slow, yet frequently provide a valid PTE even before the data address calculation. This results in a performance improvement, and reduced data address translation energy, in most cases.
@@ I Introduction
Since the processing of load (and store) instructions is governed by the addresses of the data accessed, most techniques to improve their processing, including DTLBs for address translation, are based upon the empirical properties of the addresses of the data memory accesses. However, some techniques have been proposed to optimize load instruction processing that rely on the identity of the instruction—the instruction’s program counter (PC)—and not on the data memory addresses. A prime example is memory dependence prediction (MDP), also called memory disambiguation in some commercial products, which is routinely used in high-end processors to improve the scheduling of load instructions. In MDP, pairs of load and store instructions, identified by their PCs, are predicted to be dependent, agnostic of the actual data memory addresses that they access [31, 30, 14]. These pairs are tracked in a relatively small PC-indexed table which is consulted to predict whether a load could be beneficially executed before prior stores, without the determination of the data addresses of the load and store instructions.
@@
PCAX is based on two important and novel empirical observations. First, a small percentage (and number) of static load instructions of a program account for a significant number of CDTLB misses. Second, quite often, when a dynamic instance of a static load has a CDTLB miss, it accesses the same virtual page (and the same PTE) as the previous dynamic instance. These observations suggest that a (relatively small) table of PTEs, accessed with the PC of a load instruction, a PC-Indexed Data Address Translation Table (PCAT), could provide the PTEs that would normally be obtained as a result of a DTLB miss. PCAX employs PCATs (loosely) coupled to the instruction caches. A PCAT is selectively accessed during instruction fetch and provides, with a reasonable accuracy, a needed PTE before the data address is generated, the CDTLB accessed, and a miss occurs. A PCAT is not meant to replace a CDTLB but rather augment it, by providing the PTE for a reference that likely will miss in the CDTLB.
@@ II Motivation
To motivate PC-indexed address translation, we present empirical observations on some characteristics of misses in conventional DTLBs (CDTLBs) when the memory address of the data for a load instruction is submitted to the CDTLB. For benchmarks, we consider a set of 84 traces from Qualcomm, for the ARM architecture. These traces, details of which are presented in Section IV, have been used in many recent papers [6, 12, 39, 5]. We consider a single page size of 4KB, and the miss rates, with a 64-entry CDTLB, range from 1% to 31%, with CDTLB MPKIs ranging from 1 to 53. (More on the traces in Section IV-A).
@@
Figure 1 shows the percentage of static loads that account for 90% of the: (i) accesses, and (ii) misses in a CDTLB. In the figure, the X-axis is the different benchmarks, and the Y-axis (note the log scale) is the percentage of static loads. We consider 3 CDTLB sizes: 64, 256 and 1024 entries, with 4-way associativity in each case. The larger sizes are representative of L2 or secondary TLBs. For a given CDTLB size, we only plot the cases for which the miss rate is greater than 1%.
The observation we make from the figure, which we do not believe has been documented before, is that a very small percentage of static loads are problem loads accounting for most (90%) of the misses in the CDTLB111 A related observation that has been known for a while is that a small percentage of static loads account for a large percentage of data cache misses [4, 51, 48]). This percentage of problem loads is typically an order of magnitude smaller than the percentage of static loads accounting for the references. For example, with a 64-entry CDTLB, in the median case, about 30% of the static loads account for 90% of the CDTLB accesses, but only about 4% of the static loads account for 90% of the CDTLB misses. From Figure 1 we also see that as the CDTLB size increases the percentage of static loads that are problem loads decreases  [^2] In rare cases the number of problem loads is higher for bigger CDTLBs. We address this in section IV-D. Overall, the percentage of problem loads is in the single digits.
@@
Figure 2 shows the percentage of time that a dynamic instance of a static load identified as a problem load accesses and misses on the same PTE as the last missing dynamic instance of the static (problem) load. The figure has 3 different DTLB sizes, 64, 256, and 1K entries. From the figure, we see that, for a very large percentage of the problem loads, the PTE that is accessed on a miss is the same as the PTE for the prior instance of the problem load. For example, at a median point, with a 64-entry CDTLB, about 75% of the loads that miss in the CDTLB access the same PTE as the prior dynamic instance. This percentage decreases as the CDTLB size increases. But even for the large CDTLB sizes, more than half of the problem loads access the same PTE as the prior dynamic instance.
@@
The empirical observations of Figures 1 and 2 suggest that if we were to track the last PTE for static loads that miss in the CDTLB, in a table indexed by the load PC, we could get the correct PTE for the data that a load instruction would reference, and would likely miss in the CDTLB, a fairly large percentage of the time. Further, using the load PC, this correct (and CDTLB-missing) PTE could be obtained while the load was being fetched, i.e., even before it is decoded and well before the address of the data it will reference is determined. This tracking needs to be done only for a small percentage of the static loads in a program [^3] We will consider the absolute number, rather than the percentage of problem loads later when we evaluate different structure sizes..
@@ III-A Overview
Motivated by the empirical observations of Section II, we propose PC-Indexed Address Translation (PCAX). The idea is to maintain PTEs for a select group of load instructions in a PC-Indexed Address Translation Table (PCAT). When one of these select loads is fetched, the PCAT is accessed to find a corresponding entry, the PTE from which is then provided in the expectation that it will be the PTE that will be needed when the address of the data referenced by that load is determined. The purpose of PCAX is to augment a CDTLB. For most loads, the CDTLB provides the PTE for the accessed data. For select load instructions, a (hopefully good) translation is provided by a PCAT.
@@ III-B PCAT Structure and Operation
A PCAT is a cache, with a tag array and a data array. An entry shown in Figure 3, has (i) tag bits, (ii) a virtual page number (VPN), (iii) the PTE for that virtual page, which we collectively abbreviate as VPN-PTE, and (iv) two additional bits to help with replacement decisions: a Useful bit and a STLBM bit. Additionally, each entry has a Valid bit. The tag bits and the VPN-PTE are self-explanatory. We will see the use of the other bits later.
Figure 3: PCAT Tag and Data Entry
Since access to a PCAT entry is closely tied to an instruction access—a PCAT entry won’t be accessed if the corresponding load instruction isn’t also accessed—it is natural for the PCAT structures to mirror the cache structures holding instructions. Accordingly, we propose a 2-level PCAT hierarchy, an L1 PCAT (PCAT1) coupled to an L1 instruction cache and an L2 PCAT (PCAT2) coupled to an L2 cache.
Entries are created in the PCAT1 when there is a miss in the CDTLB. The PC of the load instruction is used to determine a place (set and way) where the new entry will reside, and the tag bits populated. The PTE returned on the CDTLB miss, and the VPN of the accessed data address, are entered into the VPN-PTE bits of the entry, and the Useful bit is unset. If the entry is created as the result of a Secondary TLB (STLB) miss, the STLBM bit is set; we discuss its use below. When an entry in the PCAT1 provides a useful translation, the Useful bit gets set.
A newly-created PCAT1 entry moves to PCAT2 when it is evicted from PCAT1 thus creating a new entry in the PCAT2, eventually populating the PCAT2  [^4] An alternative is to create an entry in the PCAT2 at the same time as in PCAT1. Eviction from the PCAT1 is independent of eviction of the corresponding code cache block from the L1i cache. However, an entry is moved from PCAT2 to PCAT1 (resulting in an eviction of an entry from PCAT1) when the corresponding instruction block is moved from the L2 to the L1i cache, and not on a PCAT1 miss, as in a typical cache hierarchy.
@@
III-C Replacement in the PCATs
The PCATs can use normal cache replacement policies to determine a victim entry. However, we have found it beneficial to enhance a normal (e.g., LRU) replacement policy with some additional information. To preserve entries that have been deemed to be useful over those that have not been deemed so, we use the Useful bit to exclude an entry from a replacement decision when possible. Likewise, we have found it beneficial to try to preserve entries that were created as a result of an STLB miss, especially for the PCAT2, since they can potentially provide a translation that might otherwise have to be obtained via an STLB miss. The STLBM bit indicates such entries. The modified LRU replacement algorithm that we employ does not choose an entry with either its Useful or STLBM bit set, unless it doesn’t have a choice.
III-D Load Pipeline with PCATs
@@
Figure 4 shows the steps with relevant portions for the operation of a load in the instruction pipeline, which we describe below. The PCAT1 is accessed using the PC of a load instruction, alongside the fetch of the load instruction. However, we only want to access the PCAT1 for loads for which there is an entry in the PCAT1, and not for every load instruction. The latter is wasteful as most load PCs won’t have a matching entry in the PCAT. For this, we can maintain a Problem Load (PL) bit with an instruction in the L1i cache. However, few instructions in the L1i cache are problem loads and thus few PL bits are set. Indeed, even for cache blocks that contain a problem load, often there is only one problem load. Accordingly, we use an Access PCAT (APT) bit with a block in the L1i cache. When an L1i cache block is accessed, and its APT bit set, the PCAT1 is accessed.
@@
Though the PCAT1 is accessed alongside instruction fetch, its output—a VPN-PTE—isn’t needed until much later in the instruction processing pipeline. In particular, until the address for the data of the load instruction has been determined (and the VPN known). Accordingly, we propose the well-known two-step process for a cache access which is used to reduce access energy (at the expense of latency): access the tag array in the first step to determine if there is a hit, and the data array in the second step in case of a hit. On an access and a match, the PCAT1 provides a VPN-PTE, which is taken to the data address generation stage of the pipeline and maintained in a Recent PCAT Accesses Buffer (RPA buffer).
When the base register for a load instruction is read to start a data address calculation (or even earlier depending upon the pipeline) the VPN bits of the base register are compared with the VPN bits of the RPA buffer entries. On a match, the PTE from the RPA buffer is used if the VPN does not change in the address calculation (i.e., offset smaller than a page and no carry into VPN bits). In this way the access of the RPA buffer can be carried out in parallel with (or prior to) the address calculation. The VPN-PTE also contains the permission bits for the page (like a CDTLB) and permission checks are made at this point. If there is no match the data address is submitted to the CDTLB to obtain the corresponding PTE.
@@
The VPN-PTE obtained from the PCAT1, while intended for use by the load instruction whose PC was used to access the PCAT1, can also be used by other instructions in the data address generation stage of the pipeline while it resides in the RPA buffer, i.e., until it is replaced by a new VPN-PTE from the PCAT1. As has been observed by others, different load instructions that are close in the instruction stream, and that use the same base register, often access the same virtual page (and PTE) [7, 13]. Similarly we have observed that different static loads that are temporally close in the dynamic instruction stream access the same PTE (they generally have the same base register, with different offsets, and likely access different fields of the same data structure). However, they are not all ”contiguous” in the dynamic stream of memory referencing instructions; they are interleaved with other memory referencing instructions which access other VPNs (that hit in the CDTLB). We have found this RPA buffer, which fleetingly holds select PTEs, to service the address translations for a fair number of instructions, as we shall see in section IV-I.
III-E Address bits for PCAT access

### v2

# PC-Indexed Data Address Translation

@@ Abstract
This paper proposes a novel way to assist conventional data address translation. The approach, PC-Indexed Data Address Translation (PCAX), uses the PC of a load instruction, and not a data virtual address, to obtain the page table entry (PTE) for the data of the load instruction. PCAX is intended to be used for a small subset of the static loads in a program. We observe that: (i) a small subset of static loads is responsible for most of the misses in a data TLB (DTLB), and (ii) often a dynamic instance of a static load instruction accesses the same PTE as the last dynamic instance, and consider PCAX for this subset. With PCAX the effective miss rate of a conventional DTLB can be cut down by a factor of 2-3X, in many cases, and even more in some cases. PCAX is also beneficial in reducing the number of secondary TLB (STLB) misses. Since the PCAX tables are accessed alongside instruction fetch, they can tolerate relatively long access latencies while still frequently providing a valid PTE even before data address calculation. This yields an average performance improvement of 1.7% while reducing data address translation energy by 7% across 84 server traces.
PC-Indexed Data Address Translation
@@ I Introduction
Since the processing of load (and store) instructions is governed by the addresses of the data accessed, most techniques to improve their processing, including DTLBs for address translation, are based upon the empirical properties of the addresses of the data memory accesses. However, some techniques to optimize load instruction processing rely on the identity of the instruction, specifically the instruction’s program counter (PC), rather than the data memory addresses. A prime example is memory dependence prediction, also referred to as memory disambiguation in some commercial processors. This technique is widely used in high-end processors to improve load scheduling by predicting dependencies between load and store instructions based on their PCs, without relying on the actual data memory addresses they access [25, 12]. These pairs are tracked in a small PC-indexed table. This table is consulted to predict whether a load can be benefically executed before prior stores, without the determination of the data addresses of the load and store instructions.
@@
PCAX is based on two important and novel empirical observations. First, a small percentage (and number) of static load instructions of a program account for a significant number of CDTLB misses. Second, quite often, when a dynamic instance of a static load has a CDTLB miss, it accesses the same virtual page (and the same PTE) as the previous dynamic instance.
These observations suggest that a (relatively small) table of PTEs, accessed with the PC of a load instruction, a PC-Indexed Data Address Translation Table (PCAT), could provide the PTEs that would normally be obtained as a result of a DTLB miss. PCAX employs PCATs (loosely) coupled to the instruction caches. A PCAT is selectively accessed during instruction fetch and provides, with a reasonable accuracy, a needed PTE before the data address is generated, the CDTLB accessed, and a miss occurs. A PCAT is not meant to replace a CDTLB but rather augment it, by providing the PTE for a reference that likely will miss in the CDTLB.
@@ II Motivation
We present empirical observations to motivate PC-indexed address translation. These observations focus on the characteristics of misses in conventional DTLBs (CDTLBs), when the load data address is submitted to the CDTLB. For benchmarks, we consider a set of 84 traces from Qualcomm, for the ARM architecture. These traces, details of which are presented in Section IV, have been used in several recent papers [2, 8, 32]. We consider a single page size of 4KB, and the miss rates, with a 64-entry CDTLB, range from 1% to 31%, and the MPKIs from 1 to 53, as we shall see in section IV-D.
Figure 1 shows the percentage of static loads that account for 90% of the: (i) accesses, and (ii) misses in a CDTLB. In the figure, the X-axis is the different benchmarks, and the Y-axis (note the log scale) is the percentage of static loads. We consider 3 CDTLB sizes: 64, 256 and 1024 entries, with 4-way associativity in each case. The larger sizes are representative of L2 or secondary TLBs. For a given CDTLB size, we only plot the cases for which the miss rate is greater than 1%. By vast majority we mean 90% (of the total accesses or misses).
@@
The observation we make from the figure, which we do not believe has been documented before, is that a very small percentage of static loads are problem loads accounting for a vast majority of the misses in the CDTLB. A related, known observation that is that a small percentage of static loads account for a large percentage of data cache misses [1, 38]. Similarly, [39] observes that a few instruction accesses account for a vast majority of the instruction TLB misses. This percentage of problem loads is typically an order of magnitude smaller than the percentage of static loads accounting for the references. For example, with a 64-entry CDTLB, in the median case, about 30% of the static loads account for the vast majority of CDTLB accesses, but only about 4% of the static loads account for the vast majority of CDTLB misses. From Figure 1 we also see that as the CDTLB size increases the percentage of static loads that are problem loads decreases. In rare cases the number of problem loads is higher for bigger CDTLBs. We address this in section IV-D. Overall, the percentage of problem loads is in the single digits.
Figure 2: Loads Accessing same PTE
Figure 2 presents the percentage of all loads (both hits and misses in the CDLTB) that access the same PTE as the previous dynamic instance of the load. Observe that the percentage of loads accessing the same PTE as the previous instance is as high as 70-90% for many of the benchmarks. This result should not be surprising as many dynamic instances of a static load access proximate elements of a data structure, which are likely to be in proximate memory addresses, and thus mostly on the same page.
@@
Figure 3 presents the percentage of time that a dynamic instance of a static load identified as a problem load accesses and misses on the same PTE as the last missing dynamic instance of the static (problem) load. The figure has 3 different DTLB sizes, 64, 256, and 1K entries. Observe that for a very large percentage of the problem loads, the PTE that is accessed on a miss is the same as the PTE for the prior instance of the problem load. For example, at a median point, with a 64-entry CDTLB, about 75% of the loads that miss in the CDTLB access the same PTE as the prior dynamic instance. This percentage decreases as the CDTLB size increases. But even for the large CDTLB sizes, more than half of the problem loads access the same PTE as the prior dynamic instance.
Figure 4: L1i MPKI
@@
The empirical observations of Figures 1 and 3 suggest that if we were to track the last PTE for static loads that miss in the CDTLB, in a table indexed by the load PC, we could get the correct PTE for the data that a load instruction would reference, and would likely miss in the CDTLB, a fairly large percentage of the time. Further, using the load PC, this correct (and CDTLB-missing) PTE could be obtained while the load was being fetched, i.e., even before it is decoded and well before the address of the data it will reference is determined. This tracking needs to be done only for a small percentage of the static loads in a program (we will quantify the absolute number of problem later in section IV-D).
Figure 5: Total Data Pages Touched
Next we consider some key data page reference characteristics to motivate how the above observations could be exploited. Figure 5 presents the number of 4KB data pages touched by the different applications (note log scale on Y-axis). The data footprint is not very large. Figure 4 presents the L1i MPKI touched by the difference benchmarks for an L1 instruction cache (L1i) of size 32 KB. As we see, the benchmarks also have large code footprints [2, 8, 32]; for an L1 instruction cache (L1i) of size 32KB, there are 45 benchmarks with an L1i MPKI between 0-30, 11 benchmarks with an L1i MPKI between 30-45, 21 benchmarks with an L1i MPKI between 45-60 and 7 benchmarks with an L1i MPKI >> 60. The large code footprint results in very frequent code movement from an L2 cache to the L1i instruction cache, and coupled with the data referencing patterns, which we see next collectively results in a significant number of CDTLB misses, as we shall see in section IV-D.
Figure 6: PTE Reuse Characteristics
Figure 7: PTE Rereference Distance
Figure 6 presents the reuse characteristics of PTEs, that is, after an initial load reference that brings a PTE into the CDTLB, what percentage of reuses of the PTE (either from another dynamic instance of the same static load, or from dynamic instances of other static loads) occur within << 10, << 50, >> 100, or >> 1000 dynamic load instructions. From the figure we see that a very large percentage of references (and reuses) to a PTE occur within fewer than 50 dynamic load references and a very small percentage are for >> 100 references. Further, we have observed, data not shown, that dynamic instances of an average of 2-8 other static loads reference a PTE brought into a CDTLB by a dynamic instance of a static load. These results should also not be surprising because we expect references to different fields of an element of a data structure (from different static loads), which are likely to be on the same page, to occur close to each other.
Figure 7 presents the re-reference distance for a PTE. That is, between the time a PTE is brought into the CDTLB by a dynamic instance of a static load, how many other PTEs are referenced, on average, until the same PTE is re-referenced and brought into the CDTLB again by another dynamic instance of the same static load. Data is presented for re-reference distances of greater than 100, 200, 500, and 1000 PTEs. Many benchmarks have more than 50-60% of such missing PTE references with a re-reference distance greater than 100 and 200. Even the percentage of such misses with re-reference distances greater than 500 and 1000 is non-trivial (20-30%) for some benchmarks. Such large re-reference distances lead to higher miss rates with relatively small-sized CDTLBs.
To put it together, the server workloads we study exhibit large instruction working sets that cause frequent eviction and later re-fetch a static load instruction from the L1i as execution moves across different code regions. As a result, successive dynamic instances of the same static load may be separated by substantial intervals in the dynamic instruction stream. During these intervals, references to other data pages displace the earlier PTE from the CDTLB. When the static load is eventually re-fetched into the L1i, it often accesses the same virtual page as before, yet the translation is no longer present in the CDTLB. It is the interaction of (i) large code footprints, which increase the temporal separation between dynamic instances of a load, and (ii) structured data referencing patterns, which preserve per-PC PTE reuse.
This is what provides the basis for the scheme we present next: associate PTEs with a small subset of loads (which miss in the CDTLB) and bring a PTE associated with a load when the load is brought back into the L1i so that it can be used when the load executes without incurring a CDTLB miss. In other words, view the issue of CDTLB misses not from the viewpoint of the data reference stream, but from the viewpoint of the instruction reference stream which creates the data references that result in the CDTLB misses.
@@ III-A Overview
Motivated by the empirical observations of Section IV-M, we propose PC-Indexed Address Translation (PCAX). The idea is to maintain PTEs for a select group of load instructions in a PC-Indexed Address Translation Table (PCAT). When one of these select loads is fetched, the PCAT is accessed to find a corresponding entry. The PTE from this entry is then provided in the expectation that it will be the PTE that will be needed when the address of the data referenced by that load is determined. The purpose of PCAX is to augment a CDTLB. For most loads, the CDTLB provides the PTE for the accessed data. For select load instructions, a (hopefully good) translation is provided by a PCAT.
Figure 8: PCAT Tag and Data Entry
@@ III-B PCAT Structure and Operation
A PCAT is a cache with a tag array and a data array. Each entry, shown in Figure 8, contains: (i) tag bits, (ii) a virtual page number (VPN), (iii) the corresponding page table entry (PTE)—collectively referred to as the VPN-PTE pair, and (iv) two additional bits used for replacement decisions: a Useful bit, and a Secondary TLB (STLB) Miss bit. The STLB Miss (STLBM) bit is set when the entry is created as a result of a miss in the secondary TLB, and it is later used to guide replacement decisions. Each entry also includes a Valid bit. The tag and VPN-PTE fields are self-explanatory. We describe the usage of these bits in more detail later.
Since access to a PCAT entry is closely tied to an instruction access—a PCAT entry will not be accessed if the corresponding load instruction isn’t also accessed—it is natural for the PCAT structures to mirror the cache structures holding instructions. Accordingly, we propose a 2-level PCAT hierarchy, an L1 PCAT (PCAT1) coupled to an L1 instruction cache (L1i) and an L2 PCAT (PCAT2) coupled to an L2 cache. Like an L1i, the PCAT1 is virtually-indexed physically tagged (VIPT), though it could also be physically-indexed physically tagged (PIPT) as it’s output isn’t needed until later. Moreover, the decision to access the PCAT1 is governed by a access PCAT bit, which is obtained alongside L1i access, as we see below. Like an L2 cache the PCAT2 is PIPT.
Entries are created in the PCAT1 when a load instruction that misses in the CDTLB retires. The PC of the load instruction (or the cache block address of the cache block holding the instruction in an alternate design) is used to determine a place (set and way) where the new entry will reside, and the tag bits populated. The PTE returned on the CDTLB miss, and the VPN of the accessed data address, are entered into the VPN-PTE bits of the entry, and the Useful bit is unset.
When an entry in the PCAT1 provides a useful translation, the Useful bit gets set.
A newly-created PCAT1 entry moves to PCAT2 when it is evicted from PCAT1 thus creating a new entry in the PCAT2, eventually populating the PCAT2. An alternative is to create an entry in the PCAT2 at the same time as in PCAT1. Eviction from the PCAT1 is independent of eviction of the corresponding code cache block from the L1i. However, an entry is moved from PCAT2 to PCAT1 (resulting in an eviction of an entry from PCAT1) when the corresponding instruction block is moved from the L2 to the L1i, and not on a PCAT1 miss, as in a typical cache hierarchy.
@@
III-C Load Pipeline with PCATs
@@
Figure 9 illustrates the relevant stages involved in the operation of a load instruction within the pipeline. PCAX-specific operations are highlighted in red for clarity. The PCAT1 is accessed using the PC of a load instruction, alongside the fetch of the load instruction. However, we only want to access the PCAT1 for loads for which there is an entry in the PCAT1, and not for every load instruction. The latter is wasteful as most load PCs will not have a matching entry in the PCAT. To avoid this, we maintain a (pre-decoded) access PCAT bit with each instruction in the L1i to indicate whether the instruction PC was that of a problem load instruction and was used to create a PCAT entry. Since only a small subset of instructions typically fall into this category, these bits are rarely set. Indeed, even for cache blocks that contain a problem load, often there is only one problem load and we can use an access PCAT bit per block, rather than per instruction, in the L1i. The access PCAT bit is set for the corresponding cache block when the PCAT1 entry is created, which is when the load that misses in the CDTLB retires (as described in the Section III-B). The PCAT1 is accessed with the instruction PC (or cache block address of the cache block holding the instruction), when the corresponding access PCAT bit set.
@@
Though the PCAT1 is accessed alongside instruction fetch, its output—a VPN-PTE—is not needed until much later in the processing pipeline. In particular, until the address for the data of the load instruction has been determined (and the VPN known). To reduce access energy (at the expense of latency), we adopt the well-known two-step process for a cache access. The tag array is accessed first to check for a hit. If a hit occurs, the data array is accessed in the second step. On an access and a match, the PCAT1 provides a VPN-PTE, which is taken to the data address generation stage of the pipeline and maintained in a Recent PCAT Accesses Buffer (RPA buffer).
The RPA buffer is 1-2 entries, fully associative and is accessed using the data VPN. It acts like a nano-TLB, but with important differences: (i) it is only populated by select load instructions, those that hit in the PCAT1, and only with a VPN-PTE entry from the PCAT1, unlike a traditional nano-TLB which can be populated by any load instruction from a regular TLB, and (ii) it uses a FIFO replacement if there are multiple entries for a new entry from (i), and not a replacement scheme that is influenced by accesses. Similar to a nano-TLB, it can be accessed by any load instruction, using the VPN of the data address of the load instruction. The RPA buffer supports the same access rate as the CDTLB.
@@
The primary intent of the RPA buffer is to hold the VPN-PTE obtained from the PCAT1 by a (problem) load instruction, until the time the data address has been calculated for that (problem) load instruction, to see if the VPN-PTE is the correct one for the (problem) load instruction (the VPNs match), and whether or not the CDTLB needs to be accessed. However, an entry in the RPA buffer, while intended for use by the load instruction whose PC was used to access the PCAT1, can also be used by other instructions that were not involved in a PCAT1 access. As others have observed, different load instructions that are close in the instruction stream, and that use the same base register, often access the same virtual page (and PTE) [3, 11]. Similarly we have observed that different static loads, likely accessing different fields of a data structure (same base register and different offsets), are temporally close dynamically, and access the same PTE. However, they are not all contiguous in the dynamic stream; they are interleaved with other instructions accessing other VPNs (that hit in the CDTLB). We have found this RPA buffer, which fleetingly holds select PTEs, to service the address translations for a fair number of instructions, as we shall see in section IV-I.
Next we consider options for how and when the RPA buffer is accessed. An obvious option is to access it once the data address for a load instruction has been calculated, with the VPN bits from the calculated address. Noting, however, that the VPN-PTE obtained from the RPA buffer is just an optimization, with a fall back to the CDTLB, the buffer could also be accessed with a set of VPN bits that are highly likely to be the VPN bits of the data address. The VPN bits needed to determine a match could be some other set of bits involved in the data address calculation. Accordingly, when the base register for a load instruction is read for a data address calculation (or even earlier depending upon the pipeline) the VPN bits of the base register are compared with the VPN bits of the RPA buffer entries. On a match, the PTE from the RPA buffer is used if the VPN does not change as a result of the address calculation (i.e., offset smaller than a page and no carry into VPN bits). In this way the access of the RPA buffer can be carried out in parallel with (or prior to) the address calculation. If the address calculation results in a change in the VPN (an event that is expected to be quite infrequent), the entry is not used. The obtained VPN-PTE also contains the permission bits for the page (like a CDTLB) and permission checks are made at this point. If there is no match (or the VPN bits change as a result of address calculation) the data address is submitted to the CDTLB to obtain the corresponding PTE.
The RPA buffer has as many ports as the CDTLB. It is invalidated whenever the PCAT is invalidated; it need not be invalidated for other events that require a pipeline flush.
III-D Address bits for PCAT access